OpenStudy (anonymous):

\[\LARGE \int {9x^2-3\over x\sqrt 3+1}dx\]

OpenStudy (experimentx):

can you divide numerator with denominator ??/ and express it in terms of remainder???

OpenStudy (anonymous):

dIpanic did somthing similar in my previous integral question, but honestly I didn't get a thing!.. I can just rewrite it like: \[\LARGE \int \frac{9x^2}{3\sqrt3+1}-\frac{3}{x\sqrt3+1}dx\] and that's all !

OpenStudy (anonymous):

\[\LARGE \int \frac{9x^2}{x\sqrt3+1}-\frac{3}{x\sqrt3+1}dx\]

OpenStudy (anonymous):

How about a polynomial long division first?

OpenStudy (anonymous):

I'll be back in 10 mins

OpenStudy (experimentx):

no ... divide it, this is gong to be boring!!!

OpenStudy (experimentx):